# AI-Generated Images Introduce Invisible Relevance Bias to Text-Image Retrieval

### 23 Nov 2023 v1

@@ 5. Causes of Invisible Relevance Bias
In this section, we use the debiased model proposed in Section 4 to reversely analyze the causes of the invisible relevance bias. Specifically, we obtain the transformations made by the debiased retriever on AI-generated images by comparing the difference in image representations between the original retriever and the debiased retriever. The reverse process of these transformations can be seen as the cause of invisible relevance bias. We also design the relevant experiments to support this point of view. Details will be introduced in the following contents.
@@ 5.1. Transformations in Debiasing
To simplify the analysis process, we focus on the dual-encoder retrieval model that estimates the relevance score by dot product or cosine similarity between text and image representations. For the same text, image representation can directly affect the estimation of the relevance score, which can help us to find the most direct causes of invisible relevance bias, therefore, image representation is the main object in our analysis. Specifically, we aim to analyze the difference between image representations encoded by the original retriever and debiased retriever to get the transformations of debiasing from this difference and reverse these transformations to explore the causes of the invisible relevance bias.
@@
$\bm{r}$ $=v(I^{g};\theta),$
$\bm{r^{d}}$ $=v^{d}(I^{g};\theta^{\prime}).$
@@
$=[p_{1},p_{2},...,p_{n}]$
$=r\ominus r^{d},$
in which $\ominus$ means bitwise subtraction between $r$ and $r^{d}$. Then, we perform two-dimensional visualization of the $\bm{r}$, $\bm{r^{d}}$ and $\bm{p}$ of all images in datasets to try to find the patterns from them. The T-SNE visualization is shown in Figure 7. This result shows that compared with the scattered image representations, the transformations vector $\bm{p}$ shows an obvious aggregation phenomenon. This phenomenon shows that there is consistency in the transformations performed by the debiased retriever on AI-generated images with very different semantics. The role of the debiased retriever is stripping the transformations vector $\bm{p}$ out of the representations of AI-generated images.
Figure 7. T-SNE visualization of image representations and transformations in datasets.
@@ 5.2. Reversing the Transformations
Therefore, we conclude that an unbiased representaion $\bm{r^{d}}$ becomes the representaion $\bm{r}$ with bias because $\bm{r^{d}}$ is added bitwisely by a vector $\bm{-p}=[-p_{1},-p_{2},-p_{3},...,-p_{n}]$. Combining the phenomenon shown in Figure 7 that there is consistency in the transformations vector $\bm{p}$ on AI-generated images with very different semantics, we can get the causes of invisible relevance bias in AI-generated images: AI-generated images cause the image encoder in the retriever to embed additional information to their representations. This information has the following three characteristics:
(1) When this information is embedded into the representation outputted from the visual encoder, it can amplify the relevant information to get the higher relevance score.
(2) This information cannot be reflected in visual semantics in a visible way, but can only be extracted by neural network-based visual models.
(3) This information has a certain consistency in AI-generated images with different semantics. This information may be like the watermark that is universal information for the image generation model and can be expressed by different neural network-based visual models such as image encoder in retriever.
For the above three points, we propose experiments to support them, which will be introduced in the following contents.
@@ 5.3. Supporting Experiments
For the first point, we design a direct experiment to support it. We apply the bitwise addition of the reverse transformations vector $\bm{-p}$ to the representations of real images encoded by the original, non-debiased retriever and detect whether the bias can be eliminated. The intention for this is that if this additional information (transformations vector $\bm{-p}$ ) is indeed the cause of the AI-generated images being ranked higher, then when we add these transformations to the representation of real images, the real images will also be ranked higher, thereby eliminating the ranking difference between the real and generated images. The experimental results are shown in Table 6. It is very surprising that the ranking advantage of generated images over real images caused by invisible relevance bias is not only eliminated but reversed by simply bitwisely adding the transformations vector $\bm{-p}$ to the representation of the real images without any debiasing training. This proves that the reverse transformations vector $\bm{-p}$ we found is an important cause of the invisible relevance bias and it is implicit in the AI-generated images and can be embedded into the representations by the image encoder.
For the second point, human evaluation in Section 2.2.2 and retrieval performance in Section 2.2.1 have shown that AI-generated images do not introduce additional visual semantics compared to their real images, indicating that the additional information is invisible. Besides, the ranking bias detected in Section 3.3 and Section 3.4 shows that this additional information can be captured by the image encoder and reflected in the relevance score between the queries.
Table 6. Effect of reverse transformations vector $\bm{-p}$ on invisible relevance bias. The retrieval model is VSE trained on Flicker30k without any debiasing training.
Flicker30k+AIGC
@@ 6.1. Artificial Intelligence Generated Content
With the development of generation models, AI-generated content (AIGC) becomes more realistic and less discernible (Agnese et al., 2020; Cao et al., 2023; Wu et al., 2023). The use of AI to generate content has greatly changed the way of content generation. It reduced the cost of content generation and improved the efficiency than traditional compared to traditional content generation pipeline biased on humans (Hanley and Durumeric, 2023; Spitale et al., 2023). For example, Stable Diffusion (Yang et al., 2023) and DALL-E-3 (Betker et al., 2023) can generate high-quality images by simply following a short text description of uses’ input. However, as AIGC is gradually applied to various content production pipelines in society and the Internet, some potential risk concerns are also worthy of consideration. Some studies have found that AI-generated content has risks in many aspects, such as discrimination, privacy leakage, ethics, and safety (Deshpande et al., 2023; Zhuo et al., 2023; Chen and Shu, 2023; Jiang et al., 2023; Su et al., 2023). A recent study suggests that texts generated by AI will introduce bias in text modal information retrieval, causing the retriever to rank them higher (Dai et al., 2023). Therefore, in the context of AIGC being widely used, the potential risks it may bring should be carefully studied.
@@ 7. Conclusion
This paper delves into the exploration of source bias in text-image retrieval, focusing particularly on the phenomenon of invisible relevance bias introduced by AI-generated images. Through the construction of a comprehensive benchmark with visually similar semantics between real and AI-generated images, we unveil the prevalence of this bias in text-image retrieval models. Our findings underscore that AI-generated images tend to be ranked higher by retrieval models, despite lacking more visually relevant features to the query than real images. Regardless of the model’s training data and architecture, ranging from models trained from scratch to those pre-trained on extensive image-text pairs, and encompassing both dual-encoder and fusion-encoder models, the invisible relevance bias persists. Moreover, mixing AI-generated images into the training data exacerbates the invisible relevance bias, creating a self-perpetuating vicious cycle where these AI-generated images gain more exposure in massive datasets, increasing their likelihood of being integrated into retrieval model training, thus intensifying the bias further. In response to these challenges, we propose a novel debiasing training method to mitigate the invisible relevance bias effectively. This method can dynamically adjust the preference to AI-generated images of text-image retrieval models by adjusting the sampling probability. Then, we use our proposed debiasing method to reversely determine that the causes of invisible relevance is that the AI-generated images cause the image encoder to introduce additional information into their representation. This information has a certain consistency in different generated images and can make the retriever estimate a higher relevance score. Our proposed method and view, supported by rigorous experiments, serves as a crucial step toward understanding and rectifying the invisible relevance bias cased by AI-generated images in text-image retrieval models.

### 15 Jan 2024 v3

@@ 5. Causes of Invisible Relevance Bias
In this section, we use the debiased model proposed in Section 4 to reversely analyze the causes of the invisible relevance bias. Specifically, we discern the transformations performed by the debiased retriever on AI-generated images through a comparative analysis of the difference in image representations between the original retriever and the debiased retriever. The reverse process of these transformations can be seen as the cause of invisible relevance bias. We also design the experiments to support this point of view.
@@ 5.1. Transformations in Debiasing
To simplify the analysis process, we focus on the dual-encoder retrieval model that estimates the relevance score by computing vector similarity between text and image representations. We use image representation as the main object in our analysis because for the same text, image representation can directly affect the estimation of the relevance score, which can help us to find the most direct causes of invisible relevance bias. Specifically, we analyze the difference between image representations encoded by the original retriever and debiased retriever to get the transformations of debiasing from this difference and reverse these transformations to explore the causes of the invisible relevance bias.
@@
$\bm{r}=v(I^{g};\theta),\quad\bm{r^{d}}=v^{d}(I^{g};\theta^{\prime}).$
@@
$=[p_{1},p_{2},...,p_{n}],$
Then, we perform two-dimensional visualization of the $\bm{r}$, $\bm{r^{d}}$ and $\bm{p}$ of all images in datasets to try to find the patterns from them. The T-SNE visualization in Figure 6 shows that compared with the scattered image representations, the transformations vector $\bm{p}$ shows an obvious aggregation phenomenon. This indicates that there is consistency in the transformations performed by the debiased retriever on AI-generated images with very different semantics.
Figure 6. T-SNE visualization of image representations and transformations vector $\bm{p}$.
@@ 5.2. Reversing the Transformations
Therefore, we conclude that an unbiased representaion $\bm{r^{d}}$ becomes the representaion $\bm{r}$ with bias because $\bm{r^{d}}$ is added bitwisely by a vector $\bm{-p}=[-p_{1},-p_{2},-p_{3},...,-p_{n}]$. Combining the phenomenon shown in Figure 6 that there is consistency in the transformations vector $\bm{p}$ on AI-generated images with very different semantics, we can get the causes of invisible relevance bias in AI-generated images: AI-generated images cause the image encoder in the retriever to embed additional information to their representations. This additional information is the direct cause of invisible relevance bias and can be obtained by the difference in image representations between original and debiased retriever. This information has the following three characteristics: (1) This information cannot be reflected in a visible way, but can only be embedded by neural network-based models. (2) When this information is embedded into the representation of the image, it can amplify the query-image relevance to produce a higher score. (3) This information has a certain consistency in AI-generated images with different semantics. This information may be like the watermark that is universal information for the image generation model and can be expressed by neural network visual models such as image encoder. We design the experiments to support these three points.
@@ 5.3. Supporting Experiments
Table 6. Effect of reverse transformations vector $\bm{-p}$ on test set of Flicker30k. The retrieval model is VSE trained on Flicker30k without any debiasing training.
@@
For the first point, human evaluation in Section 2.2.2 and retrieval performance in Section 2.2.1 have shown that AI-generated images do not introduce additional visual semantics compared to their real images, indicating that the additional information is invisible. Besides, the ranking bias detected in Section 3.3 and Section 3.4 shows that this additional information can be embedded by the image encoder and produce higher relevance score than real images.
For the second point, we design a direct experiment to support it. We apply the bitwise addition of the reverse transformations vector $\bm{-p}$ to the representations of real images encoded by the original, non-debiased retriever and detect whether the bias can be eliminated. The intention for this is that if this additional information ($\bm{-p}$) is indeed the cause of the higher ranking of AI-generated images, then by incorporating this information into the representation of real images, the real images will similarly attain a higher ranking. Consequently, this would mitigate the ranking disparity between real and generated images. The experimental results are shown in Table 6. It is very surprising that the ranking advantage of generated images over real images caused by invisible relevance bias is not only eliminated but reversed by simply bitwisely adding $\bm{-p}$ to the representation of the real images without any debiasing training. This proves that the reverse transformations vector $\bm{-p}$ we found is an important cause of the invisible relevance bias. It is implicit in the AI-generated images and can be embedded into the representations by the image encoder.
@@ 6.1. Artificial Intelligence Generated Content
With the development of generation models, AI-generated content (AIGC) becomes more realistic and less discernible (Agnese et al., 2020; Cao et al., 2023; Wu et al., 2023). The use of AI to generate content has greatly changed the way of content generation. It reduced the cost of content generation and improved the efficiency than traditional compared to traditional content generation pipeline biased on humans (Hanley and Durumeric, 2023; Spitale et al., 2023). For example, Stable Diffusion (Yang et al., 2023) and DALL-E-3 (Betker et al., 2023) can generate high-quality images by simply following a short text description of uses’ input. However, as AIGC is gradually applied to various content production pipelines in society and the Internet, some potential risk concerns are also worthy of consideration. Some studies have found that AI-generated content has risks in many aspects, such as discrimination, privacy leakage, ethics, and safety (Deshpande et al., 2023; Zhuo et al., 2023; Chen and Shu, 2023; Jiang et al., 2023; Su et al., 2023). A recent study finds that texts generated by AI introduce bias in text retrieval, causing the retriever to rank them higher (Dai et al., 2023). This paper extends the study of this bias to cross-modal retrieval.
@@ 7. Conclusion
This paper delves into the exploration of the bias in text-image retrieval introduced by AI-generated images. We construct a reasonable benchmark to simulate the retrieval scenarios comprising both real images and AI-generated images. Experiments on this benchmark underscores that AI-generated images tend to be ranked higher by retrieval models, despite lacking more visually relevant semantics to the queries than real images. We define this bias as invisible relevance bias. This bias is prevalent across retrieval models with varying training data and architectures, including models trained from scratch and those pre-trained on massive image-text pairs, and encompassing dual-encoder and fusion-encoder models. Moreover, mixing AI-generated images into the training data makes the bias more serious, causing a vicious cycle where AI-generated images gain more exposure from massive data, increasing their likelihood of being mixed into retrieval model training, and exacerbating the bias further. To solve this, we propose an effective debiasing method to mitigate the invisible relevance bias. Then, we use our proposed debiasing method to reversely determine that the cause of invisible relevance is that the AI-generated images cause the image encoder to introduce additional information into their representation. This additional information can make the retriever estimate a higher relevance score. Findings in this paper reveal the potential impact of AI-generated images on text-image retrieval systems in the context of the rapid development of AIGC and have implications for further research.